in a certain office, 20 employees had perfect attendance records for the month of november. this represents 80 percent of the total office staff. how many persons are employed in this office?
Answer by rothauserc(4718)

The problem is set up with the following equation:
.8x = 20 (we solve for x, 80% of x = 20)
x = 20*10/8 = 25 are employed in the office